Character’s Name: Merlin  Player’s Name: Peter Balsiger Race:  Human  Class:  ^{Mage^{Wizard}}  AL: N  Lvl: 1  Sex: M Weight: 120 lb.  Age:  18 yrs  Height: 90 in. App.:  Normal, humble man Subrace:  Religion: Mystra  Eyes: brown  Hair: black Children: none  Siblings: unknown  Homet.: unknown cave Backg.: You certainly know, don’t you ?

STR: 12 DEX: 14 CON: 16 INT: 18 WIS: 17 CHR: 13

{Hit. \atop Prob.}  {Dmg. \atop Adj..}  {Wgt. \atop All.}  Max Press  {Op. \atop Drs.}  {B.B/ \atop L.G.}
0  0  45  140  7  4%
{Rctn. \atop Adj.}  Missile Att Adj.  {Def. \atop Adj.}
0  0  0
{HP. \atop Adj.}  {Sys. \atop Shk.}  {Res. \atop Sur.}  {Pois. \atop Sav.}  Reg.
2  95%  96%  0  Nil
No.of. Lang.  Spell Lvl.  {Lrn. \atop Sp.}  {Sp/ \atop Lvl.}  {Sp. \atop Imm.}
7  9th  85%  18  0. Ill
{Mg.D. \atop Adj.}  {Bon. \atop Sp.}  {Sp. \atop Fail.}  {Sp. \atop Imm.}
3  No  0%  No
Max. Hench.  {Loy. \atop Base}  {Rctn. \atop Adj.}
5  0  1

Movement^{Base:}^{Lgh:}^{Mod:}^{Hvy:}^{Sev:}^{Encumbrance:} ^{12}^{9}^{(46 - 69)} ^{6}^{(70 - 93)} ^{3}^{(94 - 117)} ^{1}^{(118 - 140)}^{100}

Mod. ^{Sav.Throw} ^{Base} 15 14 11 12 13^{Breath}^{Weapon}^{Paralyze}^{Poison}^{Rod,Staff}^{or Wand}^{Spells}^{Petrify}^{Polym.}

AC 4 ^{Modifiers}  ^{Armor}  ^{Weight}^{Surprised}^{Shieldless}^{Rear}  ^{5}^{5}^{5} ^{Elven chain mail}  ^{1 lb.}^{Medium shield}  ^{5 lb.}
HP 6^{Base: 12}  ^{Wounds^{Some hairs broken..}}

Weapon  ^{#Att} ^{AttAdj} ^{DAdj} ^{THAC0} ^{DmgSM} ^{DmgL}  ^{Rg S}  ^{Rg M}  ^{Rg L}  ^{Wght}  ^{Type}  ^{Size}  ^{Speed} Dagger  2  1  20 1d4 1d3  1 lb. P  S  7

Special Attacks War cry :-)  ^{Ammunition} Sling bullets

Proficiencies

Languages:

common

Weapon:

sling

Non-Weapon:

hypnotism programming

{(W/-2) \atop (I/-25)}

( - )

( - )

( - )

( - )

( - )

( - )

( - )

( - )

( - )

( - )

( - )

( - )

( - )

( - )

( - )

( - )

( - )

Abilities Writing LaTeX macros (really ?)

( - )

( - )

Gear^{Location} ^{Item}  ^{Weight} Backpack  2 lb. Picks  1 lb. Leather armor  15 lb. 9 Candles  1 lb. Clear Crystal  1 lb. Robe  1 lb. Wineskin  1 lb.  Rope (50ft)  1 lb. 3 Sheet of paper  1 lb.^{Location} ^{Item}  ^{Weight}

Water Wine Rations Feed

XP 9999^{Bonus: 10%}

Gems^{8 rubies}  ^{Coins} 220^{cp} 22^{sp}  2210 ^{ep} 121 ^{gp}  1330 ^{pp}  ^{other valuabels} 5 very valuable Art Objects

Misc. Information (Magical Items, Command Words, Small Maps etc.)

Spellbook Amulet Flying Carpet

Henchmen/Animal Companions^{Name}  ^{AC}  ^{Race/Class}  ^{HD/Lvl} ^{#AT} ^{THAC0}  ^{Dmg/Effects}  ^{Skills/Abilities}  ^{HP} Wonder Dog  5  Dog  20  3  12  1d8/1d4/1d4  Loud bark  25 Found as baby, then raised till adultry. Wonder Dog II  5  Dog  20  3  12  1d8/1d4/1d4  Loud bark  25 Found as baby, then raised till adultry. (whoaw, two dogs...what luck)^{Background:}^{Background:}^{Background:}^{Background:}^{Background:}^{Background:}^{Background:}^{Background:}